How far will a ball travel horizontally before hitting the ground, if the ball is thrown from a 300ft tower, parallel to the ground at 95mph?
Answer by Alan3354(69443)

How far will a ball travel horizontally before hitting the ground, if the ball is thrown from a 300ft tower, parallel to the ground at 95mph?
------------
1st find the time to hit the ground.
h(t) = -16t^2 + 300 is the function giving the height above gnd in terms of t in seconds.
Set it = 0
-16t^2 + 300 = 0
t^2 = 300/16 = 75/4 seconds
t = ~ 4.330 seconds.
--------------------
95 mph = 95*22/15 ft/sec = 139.333... ft/sec (horizontal speed)
s = vt (distance = speed*time)
s = 139.333*4.33
s = ~ 603.33 feet